The estimated need for converters is 500 for next year and the cost of transportation and ordering cost from supplier in India varies according to size of delivery. The cost of carrying inventions is 20% in a year for the value of average storage (calculation : storage cost for the items / value of average storage ) calculate the optimal ordering size ( EOQ= SQR(2up/cs)) when you have the following bidding offer :

  1. Ordering and transportation cots are 5000 € if the order size 500 or less
  2. Ordering and transportation cots are 8000 € if the order size 50 to 200
  3. Ordering and transportation cots are 10000 € if the order size 200 to 500

178 €

per-unit cost
Answer #1

Annual Demand = 500 per year

Holding cost = 20% of cost of item

Order quantity Ordering and transportation costs
50 or less 5000
50 to 200 8000
200 to 500 10000

Case 1:

Annual Demand = 500

Ordering cost = 5000  € per order

Carrying Cost / Holding Cost = 20 % of cost of item = 0.2 *178 = 35.6 € per unit per year

EOQ =

Q* = (2*500*5000/35.6)^0.5 = 374.76 units

Case 2:

Annual Demand = 500

Ordering cost = 8000  € per order

Carrying Cost / Holding Cost = 20 % of cost of item = 0.2 *178 = 35.6 € per unit per year

EOQ = Q* = (2*500*8000/35.6)^0.5 = 474.04 units

Case 3:

Annual Demand = 500

Ordering cost = 10000  € per order

Carrying Cost / Holding Cost = 20 % of cost of item = 0.2 *178 = 35.6 € per unit per year

EOQ = Q* = (2*500*10000/35.6)^0.5 = 529.99 units